What is one side of a square whose area is 2304.
Found 2 solutions by blwinbbbles, algebrahouse.com:
Answer by blwinbbbles(106)

Start with what you know...area is found by multipling length x width and in a square length and width are the same size.
In this case length x width = 2304
so by calling length = x, you can also say the width = x....therefore
x times x = 2304
x times x can also referred to as x^2 (x squared) so...
x^2 = 2304
take the square root of both sides...
x = 48
so each side of the square is 48 (you did not include units, so I don't know if this is feet, inches, etc.
Since all sides are equal, one side of the square is 48

Answer by algebrahouse.com(1659)

"What is one side of a square whose area is 2304."

Area of a square = s^2 {area of a square is one side squared}
s^2 = 2304 {"s" is one side}
s = 48 {took the square root of both sides}
www.algebrahouse.com